Given the data in Appendix B in the textbook, calculate Kp at 25 ∘C for the reaction
MgCO3(s)⇌MgO(s)+CO2(g)

lnK = -19.45
Answer #1

ΔG0rxn = ∑[ΔGfo(products)] - ∑[ΔGfo(reactants)]

ΔG0rxn = [-569.4 -394.4] - [-1012.1]

ΔG0rxn = 48.3 kJ/mol = 48300 J/mol

ΔG0rxn = -RTlnKp; where R = 8.3145 J/mol.K; T = 25 + 273.15 = 298.15 K

lnKp = -ΔG0rxn / RT = -48300 J/mol / (8.3145 J/mol.K x 298.15 K) = -19.4839128986

lnKp = -19.4839128986

Kp = e-19.4839128986 = 3.45337819 x 10-9

Kp at 25 0C for the reaction = 3.453 x 10-9

In scientific notation answer = 3.453E-09

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")